OpenStudy (anonymous):

Solve x2 + 18 = –9x.

OpenStudy (ivettef365):

first write the equation as x^2 + 9x + 18 = 0

OpenStudy (ivettef365):

now you can factor

OpenStudy (anonymous):

thanks

OpenStudy (nurali):

x^2+18=-9x x^2+9x+18=0 (x+6)(x+3) Either x=-3 or x=-6

OpenStudy (anonymous):

thank you so much

OpenStudy (nurali):

Anytime.